# Coding for Gaussian Two-Way Channels: Linear and Learning-Based Approaches

### v1

@@ I Introduction
In addition to channel capacity, communication reliability or error probability is another important metric in information/communication theory. Currently, some researchers are focusing on examining GTWCs in terms of communication reliability. In [6], Palacio-Baus and Devroye defined error exponents for GTWCs and showed how cooperations between the two users, or using previously received symbols into creating transmit symbols, can improve the error exponents in comparison to the non-cooperative case. In [7], Vasal suggested a dynamic programming (DP)-based methodology for encoding to improve the communication reliability for GTWCs. Despite ongoing efforts to improve the communication reliability for GTWCs, the existing works still lack in providing a specific coding method and its performance evaluation under a finite block-length regime.
To the best of our knowledge, there is no framework for designing practical codes in GTWCs, which is the main motivation behind this study. A foundational coding strategy for GTWCs is to carry out a linear processing for encoding and decoding in order to simplify the system model of GTWCs and mitigate the coding complexity. It is important to note that GTWCs can be thought of as an expanded system model of feedback-enabled Gaussian one-way channels (GOWCs), where a linear coding framework for GOWCs with feedback has been well developed. The seminal work done by Schalkwijk and Kailath in [8] introduced a simple linear encoding for GOWCs that can achieve doubly exponential decay in the probability of error upon having noiseless feedback information. In [9], Butman introduced a general framework for linear coding, in which the noise may be colored, nonstationary, and correlated in GOWCs. In [10], Chance and Love proposed a linear encoding scheme for GOWCs with noisy feedback, which is further analyzed and revealed to be an optimal structure under some conditions for linear encoding by [11]. There have been attempts to view the linear code design in feedback-enabled GOWCs as feedback stabilization in control theory [12] and optimal DP [13]. Overall, despite the availability of well-developed frameworks of linear coding for feedback-enabled GOWCs, such a linear framework has not been developed for GTWCs, which is one of the main motivations of this work. Linear processing offers the significant advantage of low complexity for encoding/decoding with a simplified system model. However, a significant limitation of linear coding is its inherent constraint on producing optimal codes because of the linearity.
@@
In this paper, we aim to bridge the gaps between the two pieces of literature on GOWCs and GTWCs. The design of coding schemes for GTWCs faces the three following challenges.
(C-1) Coupling of user encoding processes. In GTWCs, each user generates a transmit symbol at each time step using its encoder. The encoding process utilizes its own message and the previously received symbols from the other user. Here, using the receive symbols as feedback information can enhance communication reliability, as demonstrated in feedback-enabled GOWCs [8, 10, 14, 17]. The generated transmit symbols are then exchanged between the users over time, resulting in a coupling effect in their encoding processes. This coupling intertwines the encoding and decoding operations of the users, ultimately affecting the overall system behavior.
(C-2) Requirement of effective decoding to capture the coupling effect. It is crucial for the decoding process at each user to effectively capture the coupling effect introduced by the encoders. This requirement necessitates a joint design of encoders and decoders for both users, adding complexity and posing a significant challenge in the design of decoders.
(C-3) Need for efficient power management within power constraints. Power management faces significant challenges due to the coupled encoding processes and the consideration of average power constraints for each user. Furthermore, when designing encoding schemes that utilize feedback information, power control over the sequence of transmit symbols becomes crucial for achieving robust error performance, as demonstrated in feedback-enabled GOWCs [8, 10, 14, 17]. Therefore, it is essential to incorporate an effective power control strategy into the encoder design for both users.
In light of the aforementioned challenges, we propose two distinct two-way coding strategies for GTWCs: (i) linear coding, which is based on tractable mathematical derivations, and (ii) non-linear coding, which is inspired by recent advancements in deep learning. The major contributions of the paper are summarized below.
@@
  • We propose a linear coding scheme for GTWCs. We first formulate a signal model for linear coding that captures the coupling of encoders to address (C-1). We then derive an optimal form of linear decoding at each user as a function of encoding schemes of the users, which addresses (C-2). To obtain a tractable solution, we convert the sum-error minimization problem to the weighted sum-power minimization problem. In weighted sum-power minimization, we provide new insights on user cooperation by exploring the relationship between the channel noise ratio and the weight coefficient imposed in weighted sum-power minimization. Building upon the insights presented in [6] that a user with lower channel noise can act as a helper by providing feedback, our work extends this concept to address weighted sum problems. In particular, whether a user would act as a helper is affected not only by the channel ratio but also the weight coefficient. We then develop an efficient algorithm for joint encoder-decoder design under power constraints for (C-3). Lastly, we discuss how to extend our linear coding approach to medium/long block-lengths.
  • We propose a novel learning-based coding architecture via RNN to capture the three aforementioned challenges of (C-1), (C-2), and (C-3). Specifically, to address (C-1), we construct a pair of interactive RNNs for encoding, where the previous output of one user’s RNN is fed into the other user’s RNN as a current input. The interactive RNNs effectively capture user cooperation in GTWCs through a novel utilization of RNNs, inspired by the successful application for feedback utilization in GOWCs [17]. To address (C-3), we introduce a power control layer at each user’s encoder, which we prove satisfies all power constraints asymptotically. To address (C-2), we adopt bi-directional RNNs with an attention mechanism to fully exploit correlation among receive symbols, in which the encoders’ coupling behavior is implicitly captured as a form of symbols. To address all these three challenges jointly in the coding architecture, we train the overall encoders/decoders at the users via auto-encoder.
@@ II-A Transmission Model
where ${n}_{1}[k]\sim\mathcal{N}(0,\sigma_{1}^{2})$ and ${n}_{2}[k]\sim\mathcal{N}(0,\sigma_{2}^{2})$ are Gaussian noises, which are independent of each other and across the channel use index $k$. We consider an average power constraint for transmission at the users as
@@ II-B Functional Form of Encoding and Decoding
The goal of exchanging transmit symbols among the users is to transmit a message available at each user to the other. As shown in Fig. 1, we consider that each User $i$, $i\in\{1,2\}$, aims to transmit a unique message represented by a bit vector, ${\bf b}_{i}\in\{0,1\}^{K_{i}}$, to the other user, where $K_{i}$ is the number of source bits at User $i$. We first provide a general functional form of encoding and decoding at the users.
@@ II-B1 Encoding
User $i$ encodes the bit vector ${\bf b}_{i}\in\{0,1\}^{K_{i}}$ to generate the $N$ transmit symbols, $\{{x}_{i}[k]\}_{k=1}^{N}$. The coding rate at User $i$ is defined by $r_{i}=K_{i}/N$. Rather than encoding the bit vector ${\bf b}_{i}$ solely to generate the transmit symbols $\{{x}_{i}[k]\}_{k=1}^{N}$, we consider a joint encoding scheme that leverages both the bit vector and the receive symbols in GTWCs. The utilization of receive (or feedback) symbols in generating transmit symbols has been demonstrated to effectively enhance communication reliability in feedback-enabled GOWCs [8, 10, 14, 17]. This observation inspires us to utilize feedback in the framework of GTWC encoding. In particular, at time $k$, the encoding at User $i$ is described as a function of the bit vector ${\bf b}_{i}$ and the $k-1$ receive symbols $\{y_{i}[j]\}_{j=1}^{k-1}$ in (1). We denote an encoding function of User $i$ at time $k$ as $f_{i,k}:\mathbb{R}^{K_{i}+k-1}\to\mathbb{R}$. We can subsequently represent the encoding of User $i$ at time $k$ as
@@ II-C Optimization of Encoders and Decoders
Since our goal is to design practical codes with finite block-lengths, we consider error probability as our main performance indicator. To quantify the error probability, we consider two different metrics of interest. The first metric is block error rate (BLER), which is the ratio of the number of incorrect bit vectors to the total number of bit vectors. Formally, the BLER of User $i$’s bit vector is defined as $\text{BLER}_{i}=\text{Pr}[\{{\bf b}_{i}\neq\hat{\bf b}_{i}\}]$, $i\in\{1,2\}$. The other metric is bit error rate (BER), which is the ratio of the number of incorrect bits to the total number of bits in bit vectors. We first represent the $\ell$-th entry of ${\bf b}_{i}$ and $\hat{\bf b}_{i}$ as ${b}_{i}[\ell]\in\{0,1\}$ and $\hat{b}_{i}[\ell]\in\{0,1\}$, respectively. Then, the BER of ${b}_{i}[\ell]$ is defined by $\text{BER}_{i,\ell}=\text{Pr}[\{{b}_{i}[\ell]\neq\hat{b}_{i}[\ell]\}]$, $i\in\{1,2\}$ and $\ell\in\{1,...,K_{i}\}$. It is worth noting that BER values can differ for each individual bit entry, i.e., $\text{BER}_{i,\ell}\neq\text{BER}_{i,m}$ for $\ell\neq m$, depending on the design of encoding/decoding functions in (3)-(4). To achieve the goal of minimizing BERs across all entries, our objective is to minimize the average BER over all entries. Formally, we define the average BER of User $i$’s bit vector as $\text{BER}_{i}=(1/K_{i})\sum_{\ell=1}^{K_{i}}\text{BER}_{i,\ell}=(1/K_{i})\sum_{\ell=1}^{K_{i}}\text{Pr}[\{{b}_{i}[\ell]\neq\hat{b}_{i}[\ell]\}]$. In this work, we assume that all bits in ${\bf b}_{1}$ and ${\bf b}_{2}$ are independent and identically distributed (i.i.d.) to one another.
@@
where ${\bf n}_{i}=[n_{i}[1],...,n_{i}[N]]^{\top}$, $i\in\{1,2\}$. The expectation in (6) is taken over the distributions of the bit vectors, ${\bf b}_{1}$ and ${\bf b}_{2}$, and the noise terms, ${\bf n}_{1}$ and ${\bf n}_{2}$; the transmit symbol $x_{i}[k]$ in (6) is the encoding output as a function of its local bit vector ${\bf b}_{i}$ and the receive symbols, $y_{i}[1],...,y_{i}[k-1]$, shown in (3). The receive symbols, $y_{i}[1],...,y_{i}[k-1]$, are dependent on the noises ${\bf n}_{1}$ and ${\bf n}_{2}$ and the other user’s encoding process, i.e., the bit vector of the other user ${\bf b}_{\bar{i}}$, through the successive encoding operation of the both users as shown in (1) and (3).
Solving (5)-(6) is non-trivial and challenging since the encoding processes of both the users (in (3)) are coupled to each other in a causal manner, and the coupling effect should be incorporated into the design of the encoders/decoders of both users. To address these challenges, we propose two distinct two-way coding approaches. First, we propose a linear coding approach with low complexity for designing the encoders/decoders in Sec. III. Despite the low design complexity, however, the linear assumption on coding may limit the performance due to limited degrees of flexibility in the design of encoders/decoders. To allow higher degrees of freedom for coding, we propose a non-linear coding approach based on deep learning in Sec. IV.
@@ III-A Signal Model for Linear Coding
To conduct linear coding for GTWCs in Fig. 1, we consider a message symbol $m_{i}\in\mathbb{R}$, where User $i$, $i\in\{1,2\}$, converts the bit vector ${\bf b}_{i}\in\{0,1\}^{K_{i}}$ to the message symbol $m_{i}$ through $2^{K_{i}}$-ary pulse amplitude modulation (PAM)  [18]. Note that $\mathbb{E}[m_{i}]=0$ and $\mathbb{E}[m^{2}_{i}]=1$. We represent the $N$ receive symbols at User $i$ (in (1)) in vector form as ${\bf y}_{i}=[y_{i}[1],...,y_{i}[N]]^{\top}\in\mathbb{R}^{N\times 1}$. Then, we can rewrite the expressions in (1) as
@@
${\bf x}_{2}$ $={\bf g}_{2}m_{2}+{\bf F}_{2}{\bf y}_{2},$ (11)
by expressing ${\bf g}_{1}$, ${\bf F}_{1}$, ${\bf g}_{2}$, and ${\bf F}_{2}$ as functions of $\tilde{\bf g}_{1}$, $\tilde{\bf F}_{1}$, $\tilde{\bf g}_{2}$, and $\tilde{\bf F}_{2}$. The detailed derivation of the functional expressions is provided in Appendix A. Given the functional expressions, both ${\bf F}_{1}$ and ${\bf F}_{2}$ are strictly lower triangular.
@@
$\mathbb{E}\big{[}\|{\bf x}_{1}\|^{2}\big{]}$ $=\|{\bf g}_{1}\|^{2}+\|{\bf F}_{1}{\bf g}_{2}\|^{2}+\|{\bf F}_{1}{\bf F}_{2}\|_{F}^{2}\sigma_{1}^{2}+\|{\bf F}_{1}\|_{F}^{2}\sigma_{2}^{2},$ (16)
$\mathbb{E}\big{[}\|{\bf x}_{2}\|^{2}\big{]}$ $=\|({\bf I}+{\bf F}_{2}{\bf F}_{1}){\bf g}_{2}\|^{2}+\|{\bf F}_{2}{\bf g}_{1}\|^{2}+\|{\bf F}_{2}({\bf I}+{\bf F}_{1}{\bf F}_{2})\|_{F}^{2}\sigma_{1}^{2}+\|{\bf F}_{2}{\bf F}_{1}\|_{F}^{2}\sigma_{2}^{2},$ (17)

### v2

@@ I Introduction
In addition to channel capacity, communication reliability or error probability is another important metric in information/communication theory. Currently, some researchers are focusing on examining GTWCs in terms of communication reliability. In [6], Palacio-Baus and Devroye defined error exponents for GTWCs and showed how cooperations between the two users, or using previously received symbols into creating transmit symbols, can improve the error exponents in comparison to the non-cooperative case. In [7], Vasal suggested a dynamic programming (DP)-based methodology for encoding to improve the communication reliability for GTWCs. Despite ongoing efforts to improve the communication reliability for GTWCs, the existing works still lack in providing a finite block-length coding scheme and its performance evaluation.
To the best of our knowledge, there is no framework for designing practical codes in GTWCs, which is the main motivation behind this study. A foundational coding strategy for GTWCs is to carry out a linear processing for encoding and decoding in order to simplify the system model of GTWCs and mitigate the coding complexity. It is important to note that GTWCs can be thought of as an expanded system model of feedback-enabled Gaussian one-way channels (GOWCs), where a linear coding framework for GOWCs with feedback has been well developed. The seminal work done by Schalkwijk and Kailath in [8] introduced a simple linear encoding for GOWCs that can achieve doubly exponential decay in the probability of error upon having noiseless feedback information. In [9], Butman introduced a general framework for linear coding, in which the noise may be colored, nonstationary, and correlated in GOWCs. In [10], Chance and Love proposed a linear encoding scheme for GOWCs with noisy feedback, which is further analyzed and revealed to be an optimal structure under some conditions for linear encoding by [11]. There have been attempts to view the linear code design in feedback-enabled GOWCs as feedback stabilization in control theory [12] and optimal DP [13]. While the aforementioned studies focus on GOWCs with passive feedback, where the receiver feeds back its received symbol without alteration, Ben-Yishai and Shayevitz proposed a linear coding scheme for active feedback, where the receiver is allowed to employ coding for feedback [14]. Overall, despite the availability of well-developed frameworks of linear coding for feedback-enabled GOWCs, such a linear framework has not been developed for GTWCs, which is one of the main motivations of this work. Linear processing offers the significant advantage of low complexity for encoding/decoding with a simplified system model. However, a significant limitation of linear coding is its inherent constraint on producing optimal codes because of the linearity.
@@
Although GTWCs can be understood as an expanded system model of feedback-enabled GOWCs, it is not trivial to generalize the linear and learning-based coding methods of feedback-enabled GOWCs to GTWCs, since the objectives and system models are different from each other. While the objective of feedback-enabled GOWCs is to improve communication performance for the one-way transmission from the transmitter to the receiver [8, 10, 15, 18], the objective of GTWCs is to improve communication performance for the two-way transmissions from User 1 to 2 and from User 2 to 1. The performance of these two-way transmissions can be balanced through the cooperation of the users. We note that the cooperation incurs a coupling effect in their encoding processes, intertwining the encoding and decoding operations of the users. To capture this two-way cooperation and the encoder’s coupling, a novel system model is essential for GTWCs. In contrast, in feedback-enabled GOWCs, the receiver only acts as a helper while the transmitter is the recipient of help, which does not capture the two-way cooperation and the encoder’s coupling.
We specify the three challenges in the design of coding methods for GTWCs as below.
(C-1) Coupling of user encoding processes. It is important to understand how to use receive symbols in generating transmit symbols (i.e., for encoding) at both users in order to improve the user’s reliability. However, designing these encoding methods is challenging, since using the received symbols in generating transmit symbols at both users across multiple channel uses in GTWCs will incur a coupling effect in their encoding processes. This coupling intertwines the encoding and decoding operations of the users, ultimately affecting the overall system behavior.
(C-2) Requirement of effective decoding to capture the coupling effect. It is crucial for the decoding process at each user to effectively capture the coupling effect introduced by the encoders in GTWCs. This requirement necessitates a joint design of encoders and decoders for both users, adding complexity and posing a significant challenge in the design of decoders, compared to feedback-enabled GOWCs.
(C-3) Need for efficient power management within power constraints. When designing encoding schemes that utilize feedback information, power control over the sequence of transmit symbols becomes crucial for achieving robust error performance. In GTWCs, power management becomes more complicated, as compared to feedback-enalbed GOWCs, due to the coupled encoding processes requiring proper power allocation into the message and feedback transmissions at both users. Therefore, an effective power control strategy should be investigated and incorporated into the encoder design for both users under the average power constraints in GTWCs.
In this paper, we aim to bridge the gaps between the two pieces of literature on GOWCs and GTWCs by addressing the above challenges. In light of the aforementioned challenges, we propose two distinct two-way coding strategies for GTWCs: (i) linear coding, which is based on tractable mathematical derivations, and (ii) non-linear coding, which is inspired by recent advancements in deep learning.
The major contributions of the paper are summarized below.
@@
  • We propose a linear coding scheme for GTWCs. We first formulate a signal model for linear coding that captures the coupling of encoders to address (C-1). We then derive a maximum likelihood decoding, i.e., an optimal form of linear decoding, at each user as a function of encoding schemes of the users, which addresses (C-2). To obtain a tractable solution, we convert the sum-error minimization problem to the weighted sum-power minimization problem. In weighted sum-power minimization, we provide new insights on user cooperation by exploring the relationship between the channel noise ratio and the weight coefficient imposed in weighted sum-power minimization. Building upon the insights presented in [6] that a user with lower channel noise can act as a helper by providing feedback, our work extends this concept to address weighted sum problems. In particular, whether a user would act as a helper is affected not only by the channel ratio but also by the weight coefficient. We then develop an efficient algorithm for joint encoder-decoder design under power constraints for (C-3). Lastly, we discuss how to extend our linear coding approach to medium/long block-lengths.
  • We propose a novel learning-based coding architecture to capture the three aforementioned challenges of (C-1), (C-2), and (C-3). Specifically, to address (C-1), we introduce state propagation-based encoding to understand the encoding behavior in two-way transmissions and propose employing a pair of interactive RNNs for encoding, where the previous output of one user’s RNN is fed into the other user’s RNN as a current input. This innovative use of RNNs in an interactive manner effectively captures user cooperation across many channel uses in GTWCs, while in feedback-enabled GOWCs, only one-way encoding RNN has been designed at the transmitter [15, 18], which does not capture the two-way transmissions and the encoder’s coupling. To address (C-3), we introduce a power control layer at each user’s encoder, which we prove satisfies all power constraints asymptotically. To address (C-2), we adopt bi-directional RNNs with an attention mechanism to fully exploit correlation among receive symbols, in which the encoders’ coupling behavior is implicitly captured as a form of symbols. To address all these three challenges jointly in the coding architecture, we train the overall encoders/decoders at the users via auto-encoder.
@@ II-A Transmission Model
where ${n}_{1}[k]\sim\mathcal{N}(0,\sigma_{1}^{2})$ and ${n}_{2}[k]\sim\mathcal{N}(0,\sigma_{2}^{2})$ are Gaussian random variables representing the additive noises, which are independent of each other and across the channel use index $k$. We consider an average power constraint for transmission at the users as
@@ II-B Functional Form of Encoding and Decoding
The goal of exchanging transmit symbols among the users is to transmit a message available at each user to the other. As shown in Fig. 1, we consider that each User $i$, $i\in\{1,2\}$, aims to transmit a unique message represented by a bit vector, ${\bf b}_{i}\in\{0,1\}^{K_{i}}$, to the other user, where $K_{i}$ is the number of source bits at User $i$. Note that ${\bf b}_{i}$ denotes a random variable. In this work, we assume that all bits in ${\bf b}_{1}$ and ${\bf b}_{2}$ are independent and identically distributed (IID). We first provide a general functional form of encoding and decoding at the users.
@@ II-B1 Encoding
User $i$ encodes the bit vector ${\bf b}_{i}\in\{0,1\}^{K_{i}}$ to generate the $N$ transmit symbols, $\{{x}_{i}[k]\}_{k=1}^{N}$. The rate at User $i$ is defined by $r_{i}=K_{i}/N$. Rather than encoding the bit vector ${\bf b}_{i}$ solely to generate the transmit symbols $\{{x}_{i}[k]\}_{k=1}^{N}$, we consider a joint encoding scheme that leverages both the bit vector and the receive symbols in GTWCs. The utilization of receive (or feedback) symbols in generating transmit symbols has been demonstrated to effectively enhance communication reliability in feedback-enabled GOWCs [8, 10, 15, 18]. This observation inspires us to utilize feedback in the framework of GTWC encoding. In particular, at time $k$, the encoding at User $i$ is described as a function of the bit vector ${\bf b}_{i}$ and the $k-1$ receive symbols $\{y_{i}[j]\}_{j=1}^{k-1}$ in (1). We denote an encoding function of User $i$ at time $k$ as $f_{i,k}:\mathbb{R}^{K_{i}+k-1}\to\mathbb{R}$. We can subsequently represent the encoding of User $i$ at time $k$ as
@@ II-C Optimization of Encoders and Decoders
Since our goal is to design practical codes with finite block-lengths, we consider error probability as our main performance indicator. We consider two different metrics of interest. The first metric is the block error rate (BLER), which represents the probability that the original bit vector from one user is incorrectly recovered by the other user. Formally, the BLER of User $i$’s bit vector is defined as $\text{BLER}_{i}=\text{Pr}[\{{\bf b}_{i}\neq\hat{\bf b}_{i}\}]$, $i\in\{1,2\}$. The other metric is the bit error rate (BER), which represents the probability that each bit in the original bit vector from one user is incorrectly recovered by the other user. We first represent the $\ell$-th entry of ${\bf b}_{i}$ and $\hat{\bf b}_{i}$ as ${b}_{i}[\ell]\in\{0,1\}$ and $\hat{b}_{i}[\ell]\in\{0,1\}$, respectively. Then, the BER of ${b}_{i}[\ell]$ is defined by $\text{BER}_{i,\ell}=\text{Pr}[\{{b}_{i}[\ell]\neq\hat{b}_{i}[\ell]\}]$, $i\in\{1,2\}$ and $\ell\in\{1,...,K_{i}\}$. It is worth noting that BER values can differ for each individual bit entry, i.e., $\text{BER}_{i,\ell}\neq\text{BER}_{i,m}$ for $\ell\neq m$, depending on the design of encoding/decoding functions in (3)-(4). To achieve the goal of minimizing BERs across all entries, our objective is to minimize the average BER over all entries. Formally, we define the average BER of User $i$’s bit vector as $\text{BER}_{i}=(1/K_{i})\sum_{\ell=1}^{K_{i}}\text{BER}_{i,\ell}=(1/K_{i})\sum_{\ell=1}^{K_{i}}\text{Pr}[\{{b}_{i}[\ell]\neq\hat{b}_{i}[\ell]\}]$.
@@
where ${\bf n}_{i}=[n_{i}[1],...,n_{i}[N]]^{\top}$, $i\in\{1,2\}$. The expectation in (6) is taken over the distributions of the random variables, i.e., the bit vectors, ${\bf b}_{1}$ and ${\bf b}_{2}$, and the noise terms, ${\bf n}_{1}$ and ${\bf n}_{2}$; the transmit symbol $x_{i}[k]$ in (6) is the encoding output as a function of its local bit vector ${\bf b}_{i}$ and the receive symbols, $y_{i}[1],...,y_{i}[k-1]$, shown in (3). The receive symbols, $y_{i}[1],...,y_{i}[k-1]$, are dependent on the noises ${\bf n}_{1}$ and ${\bf n}_{2}$ and the other user’s encoding process, i.e., the bit vector of the other user ${\bf b}_{\bar{i}}$, through the successive encoding operation of the both users as shown in (1) and (3).
Solving (5)-(6) is non-trivial and challenging since the encoding processes of both the users (in (3)) are coupled to each other in a causal manner, and the coupling effect should be incorporated into the design of the encoders/decoders of both users. To address these challenges, we propose two distinct two-way coding approaches. First, we propose a linear coding approach with low complexity for designing the encoders/decoders in Sec. III. Despite the low design complexity, the linear assumption on coding may limit the performance due to limited degrees of flexibility in the design of encoders/decoders. To allow higher degrees of freedom for coding, we propose a non-linear coding approach based on deep learning in Sec. IV.
@@ III-A Signal Model for Linear Coding
To conduct linear coding for GTWCs in Fig. 1, we consider a message symbol $m_{i}\in\mathbb{R}$, where User $i$, $i\in\{1,2\}$, converts the bit vector ${\bf b}_{i}\in\{0,1\}^{K_{i}}$ to the message symbol $m_{i}$ through $2^{K_{i}}$-ary pulse amplitude modulation (PAM)  [19]. Note that $m_{i}$ is a random variable. We assume that each message $m_{i}$ is chosen from a set of possible messages $\mathcal{M}_{i}$ according to a uniform distribution, given that all bits in ${\bf b}_{1}$ and ${\bf b}_{2}$ are assumed to be IID to one another. We note that the cardinality of $\mathcal{M}_{i}$ is $2^{K_{i}}$. We consider $\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{i}}}}}[m_{i}]=0$ and $\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{i}}}}}[m^{2}_{i}]=1$. We focus on the scenario, where Users 1 and 2 exchange a single pair of messages, i.e., User 1 conveys $m_{1}$ to User 2, and User 2 conveys $m_{2}$ to User 1.
We represent the $N$ receive symbols at User $i$ (in (1)) in vector form as ${\bf y}_{i}=[y_{i}[1],...,y_{i}[N]]^{\top}\in\mathbb{R}^{N\times 1}$. Then, we can rewrite the expressions in (1) as
@@
Remark 1.
There are various ways to represent the transmit signals, these representations can be converted into one another through variable changes due to linearity. Our signal model in (8)-(9) is one such representation. Other possible linear signal models and their conversions are discussed in detail in Appendix A. Among various choices of general linear signal models, we chose (8)-(9) because this form does not result in any inverse matrix appearing in the power expressions, as shown in (16)-(17). However, other linear signal models would lead to the introduction of an inverse matrix form in the power expressions, which are difficult to handle in optimization. The derivation of the power expressions when using other linear signal models is discussed in Appendix A.
@@
${\bf x}_{2}$ $={\bf g}_{2}m_{2}+{\bf F}_{2}{\bf y}_{2}.$ (11)
Here, we introduce another set of parameters, $({\bf g}_{1},{\bf F}_{1},{\bf g}_{2},{\bf F}_{2})$, as functions of $(\tilde{\bf g}_{1},\tilde{\bf F}_{1},\tilde{\bf g}_{2},\tilde{\bf F}_{2})$ to represent (10)-(11). The introduction of the new set of parameters makes the subsequent problem formulation and optimization easier to deal with. The detailed derivation of the functional expressions for $({\bf g}_{1},{\bf F}_{1},{\bf g}_{2},{\bf F}_{2})$ from $(\tilde{\bf g}_{1},\tilde{\bf F}_{1},\tilde{\bf g}_{2},\tilde{\bf F}_{2})$ is presented in Appendix B, whereas the derivation for the reverse case is provided in Appendix C. Given the functional expressions, both ${\bf F}_{1}$ and ${\bf F}_{2}$ are strictly lower triangular.
@@
$\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]}$ $=\|{\bf g}_{1}\|^{2}+\|{\bf F}_{1}{\bf g}_{2}\|^{2}+\|{\bf F}_{1}{\bf F}_{2}\|_{F}^{2}\sigma_{1}^{2}+\|{\bf F}_{1}\|_{F}^{2}\sigma_{2}^{2},$ (16)
$\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{2}\|^{2}\big{]}$ $=\|({\bf I}+{\bf F}_{2}{\bf F}_{1}){\bf g}_{2}\|^{2}+\|{\bf F}_{2}{\bf g}_{1}\|^{2}+\|{\bf F}_{2}({\bf I}+{\bf F}_{1}{\bf F}_{2})\|_{F}^{2}\sigma_{1}^{2}+\|{\bf F}_{2}{\bf F}_{1}\|_{F}^{2}\sigma_{2}^{2},$ (17)